# COADVISE: Covariate Adjustment with Variable Selection in Randomized Controlled Trials

### v4

@@ 1.1 Background
[…]
@@ 1.2 Related work
Despite its strengths, ANCOVA is not universally superior to the simple estimator. Its efficiency and credibility can diminish when the conditional outcome models for treatment and control groups differ significantly (Freedman, 2008). To address these limitations, Lin (2013) proposed including interactions between the treatment and covariates in the model. This method, known as analysis of heterogeneous covariance (ANHECOVA) (Cassel et al., 1976; Yang and Tsiatis, 2001; Ye et al., 2023b), allows the regression model to capture heterogeneity in the treatment effect across covariates. Although ANHECOVA has historical roots, Lin (2013) was instrumental in establishing its advantages through a design-based lens. This estimator is not only unbiased but also asymptotically more efficient than the simple estimator under a wide range of conditions, without relying on restrictive model assumptions (Lin, 2013). The adoption of ANHECOVA has expanded in recent years, particularly for its ability to enhance efficiency in RCTs employing covariate-adaptive randomization schemes (Bannick et al., 2023). The U.S. FDA has endorsed its use, especially in evaluating conditional treatment effects, further strengthening its role in modern controlled trials (FDA, 2023; Van Lancker et al., 2023). Recent work by Reluga et al. (2024) provided a unified perspective on the relationships between the simple estimator, ANCOVA, and ANHECOVA, providing conditions under which ANHECOVA uniformly dominates other methods in terms of efficiency and precision.
@@
Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009; Kang et al., 2007). The AIPW estimator combines outcome modeling with inverse treatment probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities. This dual-model approach allows AIPW to achieve semiparametric efficiency—that is, to attain the semiparametric variance lower bound—when both the treatment and outcome models are correctly specified. In the setting of RCTs considered in this paper, the treatment assignment mechanism is known by design, so only correct specification of the outcome model is required. As a result, AIPW can be asymptotically most efficient among regular estimators of the ATE in this context. Moreover, AIPW remains consistent even when the outcome model is misspecified, owing to the robustness conferred by randomization. This robustness, combined with its efficiency, makes AIPW particularly attractive in practice. Recent work has extended AIPW methodology to accommodate more complex randomization schemes. For example, Bannick et al. (2023) consider covariate-adaptive randomization, in which treatment assignment probabilities are stratified and potentially correlated with baseline covariates. In certain structured settings, such as when allocation probabilities are constant within strata, AIPW can still achieve semiparametric efficiency with only a correctly specified outcome model. However, under more general covariate-adaptive designs, correct modeling of the treatment assignment mechanism becomes necessary, mirroring challenges found in observational studies. Additionally, Ye et al. (2023a, b) propose robust variance estimators for AIPW under multi-arm RCT settings, which we adopt in our implementation (see Section 2.4).
@@
As mentioned in Section 1.1, all covariate adjustment methods encounter difficulties in ensuring parameter convergence when faced with high-dimensional covariate data. As a result, a key preliminary step is variable selection—ideally choosing a set of variables smaller than the sample size $N$—that are most predictive of the outcome. Cho and Yang (2023) and Belloni et al. (2014) proposed strategies tailored to covariate adjustment in observational studies, and Yang et al. (2020) presented complementary methods in the context of data integration. Following these prior works, we aim to select covariates that strongly predict the outcome.
@@
For binary or categorical outcomes, research on covariate adjustment has also made significant progress (Ye et al., 2023a; Moore and van der Laan, 2009; Jiang et al., 2017; Guo and Basse, 2023; Cohen and Fogarty, 2024), though variable selection has not been a primary focus in these studies. Notably, Van Lancker et al. (2024) introduced a unified framework for variable selection combined with the AIPW estimator. Their approach also leverages double/debiased machine learning (DML) algorithms (Chernozhukov et al., 2018), employing techniques such as sample splitting and cross-fitting for model fitting and outcome prediction during covariate adjustment. This integration of machine learning with semiparametric methods represents a promising direction for addressing high-dimensional covariates while maintaining the robustness and efficiency of AIPW-based estimators.
@@ 1.3 Our contributions
AIPW (Bannick et al., 2023) Delta method $\checkmark$ $\checkmark$ $\checkmark$ if (i) or (ii) in Theorem 1 holds
@@ 2.1 Notation and set-up
We consider a binary treatment, denoted by $A\in\{0,1\}$, where 111 represents the treated group and 0 represents the control group. Each participant in the sample is characterized by a baseline covariate vector $\mathbf{X}$, such as clinical characteristics. We adopt the Neyman-Rubin potential outcomes framework (Neyman, 1923; Imbens and Rubin, 2015), in which each participant has two potential outcomes, $Y(0)$ and $Y(1)$, representing the potential outcomes under control and treatment, respectively.
@@ 2.2 Causal parameter, identification assumptions, and the unadjusted estimator
where the expectation ${}\mathbb{E}\{\cdot\}$ is taken over the target population represented by the trial sample. However, we can only observe one outcome for each participant, denoted by $Y$. Throughout, we assume that (i) the observed outcome is consistent with the potential outcome under the received treatment, i.e., $Y=AY(1)+(1-A)Y(0)$; and (ii) there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]). In addition, we make the following standard identification assumptions for RCT data.
@@ 2.4 Non-linear covariate adjustment
The AIPW estimator for RCT data is given by (Bannick et al., 2023; Glynn and Quinn, 2010):
@@
A robust variance estimator for the AIPW estimator can be derived using the Delta method (Taylor’s expansion) (Ye et al., 2023a). Under certain regularity conditions (Bannick et al., 2023) and denoting $\boldsymbol{\theta}=(\theta_{0},\theta_{1})^{\prime}$, it can be shown that:
@@
The asymptotic variance for $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}=c^{\prime}\widehat{\boldsymbol{\theta}}:=f(\widehat{\boldsymbol{\theta}})$ is given by $\{\nabla f(\boldsymbol{\theta})\}^{\prime}\mathbf{V}\{\nabla f(\boldsymbol{\theta})\}=v_{00}-2v_{01}+v_{11}$, where $c=(1\ \ -1)^{\prime}$. Therefore, a consistent variance estimator is:
@@ 3.1 The COADVISE workflow
Next, we fit two parametric outcome regression models—one for $Y(1)$ and one for $Y(0)$—using a chosen link function $g(\cdot)$ from the GLM family (for example, logistic regression for a binary outcome).
Once these models are estimated, each model is applied to the data from the opposite treatment group to predict the missing potential outcomes. In other words, a model $\widehat{\mu}_{a}(\mathbf{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group $A=1-a$, for $a=0,1$. This procedure yields all the information necessary to compute the AIPW estimator in (5). Finally, we quantify uncertainty by employing the robust variance estimator in (6) (Ye et al., 2023a).
@@ 3.2 Theoretical results
  • (i) Both conditional outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, after variable selection, are specified by linear models with OLS used for estimating the regression coefficients.
  • (ii) If the outcome model for $Y(a)$ ($a=0,1$), after variable selection, is specified by a nonlinear GLM with link function $g$, then there exists a $\boldsymbol{\beta}_{a}$ as the probability limit of MLE $\widehat{\boldsymbol{\beta}}_{a}$ such that with probability 1: (a) the sign of $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ matches that of $\mathbb{E}(Y(a)\mid\mathbf{X})$; and (b) $|g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})|\leq 2|\mathbb{E}(Y(a)\mid\mathbf{X})|$.
@@ 3.3 Post-selection inference
Although Theorem 1 guarantees efficiency gain regardless of the choice of variable selection method, the uncertainty introduced by variable selection must be accounted for when conducting inference with the final AIPW estimator. This challenge is commonly referred to as the post-selection inference problem. In this section, we outline the theoretical conditions under which a variable selection method enables valid inference using the variance estimator applied to the AIPW estimator constructed from the selected covariates. Design-based theoretical results related to Lasso-based variable selection, in conjunction with the ANHECOVA estimator, have been studied in Bloniarz et al. (2016).
@@ 3.4.2 Covariate-adaptive randomization
Our second extension centers around the covariate-adaptive randomization (Bugni et al., 2018; Ye et al., 2022; Rafi, 2023; Bannick et al., 2023) including stratified permuted block randomization (Zelen, 1974) and Pocock and Simon’s minimization (Pocock and Simon, 1975; Taves, 1974) as special cases. In general, we let $\mathbf{Z}$ be a discrete baseline covariate vector used in covariate-adaptive randomization, where $\mathbf{Z}$ has $L<\infty$ distinct joint levels $\mathbf{z}_{1},\dots,\mathbf{z}_{L}$. Also, $\mathbf{Z}$ and $\mathbf{X}$ can share some common components. For example, the gender can be a factor for stratified randomization at the design stage, but it can also be incorporated into the covariate adjustment in the analysis stage. Following Bannick et al. (2023), we impose the following mild assumption on treatment assignment in covariate-adaptive RCTs. We use $\pi_{a}(\mathbf{z})=P(A=a\mid\mathbf{Z}=\mathbf{z})$ to denote the treatment assignment probability for arm $a$ within stratum $\mathbf{z}$.
@@
Furthermore, $\mathbf{\Omega}(\mathbf{z})=\text{diag}\left\{\pi_{0}(\mathbf{z}),\pi_{1}(\mathbf{z})\right\}-\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)^{\prime}\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)$ is a $\times 2$ 2 covariance matrix determined by the covariate-adaptive randomization design. In contrast, $\mathbf{\Omega}_{\text{SR}}=\text{diag}\{\pi_{0},\pi_{1}\}-(\pi_{0}~~\pi_{1})^{\prime}(\pi_{0}~~\pi_{1})$ denotes the corresponding matrix under simple randomization.
@@ 4 Numerical Experiments
[…]
@@ 4.3 Performance criteria
  • Power%: The proportion of replications where the p-value is 0.05$\leq$ 0.05≤ 0.05 for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
With $M=500$ simulations, a CPis considered significantly different from the nominal 95% coverage level if it falls outside the range $95\%\pm 1.96\times\sqrt{95\%\times 5\%/500}=[93\%,97\%]$.

### v5

@@ 1.1 Background
Despite these regulatory endorsements, significant challenges remain due to high-dimensional covariates and nonlinear relationships. These complexities necessitate robust and comprehensive analytical strategies. However, the comparative advantages and limitations of existing methods remain insufficiently explored, leaving practitioners with limited guidance and uncertainty in selecting the most appropriate approaches. For instance, in the case study presented in Section 5, the Best Apnea Interventions for Research (BestAIR) trial includes data from $N=169$ participants and $p=114$ covariates, making $p$ comparable to $N$. When adjusting for covariates using data from a single treatment arm—where $N_{1}=83$ (treated) and $N_{0}=86$ (controls)—$p$ exceeds the sample sizes. This high-dimensional setting can lead to biased and imprecise covariate-adjusted average treatment effect (ATE) estimates, as the sample size may be insufficient for reliable parameter estimation. Nevertheless, prior studies, such as Gao et al. (2024b), have identified baseline covariate imbalances in the BestAIR trial data, underscoring the need for efficient covariate-adjusted methods. This motivates our approach to improving estimation precision while addressing challenges aforementioned.
@@ 1.2 Related work
Despite its strengths, ANCOVA is not universally superior to the simple estimator. Its efficiency and credibility can diminish when the conditional outcome models for treatment and control groups differ significantly (Freedman, 2008). To address these limitations, Lin (2013) proposed including interactions between the treatment and covariates in the model. This method, known as analysis of heterogeneous covariance (ANHECOVA) (Cassel et al., 1976; Yang and Tsiatis, 2001; Ye et al., 2023b), allows the regression model to capture heterogeneity in the treatment effect across covariates. Although ANHECOVA has historical roots, Lin (2013) was instrumental in establishing its advantages through a design-based lens. This estimator is not only unbiased but also asymptotically more efficient than the simple estimator under a wide range of conditions, without relying on restrictive model assumptions (Lin, 2013). The adoption of ANHECOVA has expanded in recent years, particularly for its ability to enhance efficiency in RCTs employing covariate-adaptive randomization schemes (Bannick et al., 2025). The U.S. FDA has endorsed its use, especially in evaluating conditional treatment effects, further strengthening its role in modern controlled trials (FDA, 2023; Van Lancker et al., 2023). Recent work by Reluga et al. (2024) provided a unified perspective on the relationships between the simple estimator, ANCOVA, and ANHECOVA, providing conditions under which ANHECOVA uniformly dominates other methods in terms of efficiency and precision.
@@
Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009; Kang et al., 2007; Yuan et al., 2012; Zhang et al., 2008; Tsiatis et al., 2008). The AIPW estimator combines outcome modeling with inverse treatment probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities. This dual-model approach allows AIPW to achieve semiparametric efficiency—that is, to attain the semiparametric variance lower bound—when both the treatment and outcome models are correctly specified. In the setting of RCTs considered in this paper, the treatment assignment mechanism is known by design, so only correct specification of the outcome model is required. As a result, AIPW can be asymptotically most efficient among regular estimators of the ATE in this context. Moreover, AIPW remains consistent even when the outcome model is misspecified, owing to the robustness conferred by randomization. This robustness, combined with its efficiency, makes AIPW particularly attractive in practice. Recent work has extended AIPW methodology to accommodate more complex randomization schemes. For example, Bannick et al. (2025) consider covariate-adaptive randomization, in which treatment assignment probabilities are stratified and potentially correlated with baseline covariates. In certain structured settings, such as when allocation probabilities are constant within strata, AIPW can still achieve semiparametric efficiency with only a correctly specified outcome model. However, under more general covariate-adaptive designs, correct modeling of the treatment assignment mechanism becomes necessary, mirroring challenges found in observational studies. Additionally, Ye et al. (2023a, b) propose robust variance estimators for AIPW under multi-arm RCT settings, which we adopt in our implementation (see Section 2.4).
@@
As mentioned in Section 1.1, all covariate adjustment methods encounter difficulties in ensuring parameter convergence when faced with high-dimensional covariate data. As a result, a key preliminary step is variable selection—ideally choosing a set of variables smaller than the sample size NN—that are most predictive of the outcome. Cho and Yang (2023) and Belloni et al. (2014) proposed strategies tailored to covariate adjustment in observational studies, and Yang et al. (2020) presented complementary methods in the context of data integration. Following these prior works, we aim to select covariates that strongly predict the outcome.
@@
For binary or categorical outcomes, research on covariate adjustment has also made significant progress (Ye et al., 2023a; Moore and van der Laan, 2009; Jiang et al., 2017; Guo and Basse, 2023; Cohen and Fogarty, 2024), though variable selection has not been a primary focus in these studies. Notably, Van Lancker et al. (2024) introduced a unified framework for variable selection combined with the AIPW estimator. Their approach also leverages double/debiased machine learning algorithms (Chernozhukov et al., 2018), employing techniques such as sample splitting and cross-fitting for model fitting and outcome prediction during covariate adjustment. This integration of machine learning with semiparametric methods represents a promising direction for addressing high-dimensional covariates while maintaining the robustness and efficiency of AIPW-based estimators.
@@ 1.3 Our contributions
AIPW (Bannick et al., 2025) Delta method $\checkmark$ $\checkmark$ $\checkmark$ if (i) or (ii) in Theorem 1 holds
@@ 2.1 Notation and set-up
We consider a binary treatment, denoted by $A\in\{0,1\}$, where $1$ represents the treated group and 0 represents the control group. Each participant in the sample is characterized by a baseline covariate vector $\mathbf{X}$, such as clinical characteristics. We adopt the Neyman-Rubin potential outcomes framework (Neyman, 1923; Imbens and Rubin, 2015), in which each participant has two potential outcomes, $Y(0)$ and $Y(1)$, representing the potential outcomes under control and treatment, respectively.
@@ 2.2 Causal parameter, identification assumptions, and the unadjusted estimator
where the expectation $\mathbb{E}\{\cdot\}$ is taken over the target population represented by the trial sample. However, we can only observe one outcome for each participant, denoted by $Y$. Throughout, we assume that (i) the observed outcome is consistent with the potential outcome under the received treatment, i.e., Y=AY(1)+(1A)Y(0)Y=AY(1)+(1-A)Y(0); and (ii) there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]). In addition, we make the following standard identification assumptions for RCT data.
@@ 2.4 Non-linear covariate adjustment
The AIPW estimator for RCT data is given by (Bannick et al., 2025; Glynn and Quinn, 2010):
@@
A robust variance estimator for the AIPW estimator can be derived using the Delta method (Taylor’s expansion) (Ye et al., 2023a). Under certain regularity conditions (Bannick et al., 2025) and denoting $\boldsymbol{\theta}=(\theta_{0},\theta_{1})^{\prime}$, it can be shown that:
@@
[…]
@@ 3.1 The COADVISE workflow
Next, we fit two parametric outcome regression models—one for $Y(1)$ and one for Y(0)Y(0)—using a chosen link function $g(\cdot)$ from the GLM family (for example, logistic regression for a binary outcome).
Once these models are estimated, each model is applied to the data from the opposite treatment group to predict the missing potential outcomes. In other words, a model $\widehat{\mu}_{a}(\mathbf{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group A=1aA=1-a, for $a=0,1$. This procedure yields all the information necessary to compute the AIPW estimator in (5). Finally, we quantify uncertainty by employing the robust variance estimator in (6) (Ye et al., 2023a).
@@ 3.2 Theoretical results
  1. (i)
    Both conditional outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, after variable selection, are specified by linear models with OLS used for estimating the regression coefficients.
  2. (ii)
    If the outcome model for $Y(a)$ ($a=0,1$), after variable selection, is specified by a nonlinear GLM with link function $g$, then there exists a $\boldsymbol{\beta}_{a}$ as the probability limit of MLE $\widehat{\boldsymbol{\beta}}_{a}$ such that with probability 1:
    1. (a)
      The sign of $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ matches that of $\mathbb{E}(Y(a)\mid\mathbf{X})$; and
    2. (b)
      $|g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})|\leq 2|\mathbb{E}(Y(a)\mid\mathbf{X})|$.
@@
Remark 3 (Linear calibration for nonlinear outcome models).
When outcome models are specified via nonlinear GLMs, Condition (ii) for efficiency gains in Theorem 1 may fail, because it requires the model-induced bias to be sufficiently small—a condition that is typically untestable in practice. While Condition (i) avoids this bias requirement by using linear regressions with OLS coefficient estimates, such an approach is not always optimal; for example, nonlinear GLMs are often more natural for binary or categorical outcomes. To address this limitation, we adopt a linear calibration step motivated by recent literature (Bannick et al., 2025; Cohen and Fogarty, 2024).
Specifically, after obtaining the fitted outcome predictions $\widehat{\mu}_{1}(\mathbf{X})$ and $\widehat{\mu}_{0}(\mathbf{X})$ from nonlinear GLMs, we perform an additional calibration regression within each treatment arm. For treated participants $(A=1)$, we fit
$\texttt{lm}(Y\sim\widehat{\mu}_{1}(\mathbf{X})+\widehat{\mu}_{0}(\mathbf{X})),$
using the OLS estimates for regression coefficients, and we do the same for control participants $(A=0)$. Then, let $\widehat{\mu}_{0}^{\text{update}}(\mathbf{X})$ and $\widehat{\mu}_{1}^{\text{update}}(\mathbf{X})$ denote the predicted values for potential outcomes $Y(0)$ and $Y(1)$, respectively, from these regressions. We then replace the original $\widehat{\mu}_{a}(\mathbf{X})$ by $\widehat{\mu}_{a}^{\text{update}}(\mathbf{X})$ in the AIPW estimator $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}$, where for $a=0,1$, we update
$\widehat{\theta}_{a}=\frac{1}{N_{a}}\sum_{i=1}^{N}I(A_{i}=a)[Y_{i}-\widehat{\mu}_{a}^{\text{update}}(\mathbf{X}_{i})]+\frac{1}{N}\sum_{i=1}^{N}\widehat{\mu}_{a}^{\text{update}}(\mathbf{X}_{i}),\text{ with }N_{a}=\sum_{i=1}^{N}I(A_{i}=a).$
This correction aligns the predicted outcomes with the observed data within each treatment arm while retaining the flexibility of nonlinear GLMs, thereby ensuring that the AIPW estimator achieves efficiency gains without relying on the bias requirement in Theorem 1.
We allow users to decide whether to apply linear calibration in practice by including the option lin.cal = TRUE or FALSE in the Coadvise() function of our Coadvise R package. When users are uncertain about whether their postulated nonlinear GLM models may introduce substantial bias, enabling this additional step is recommended.
@@ 3.3 Post-selection inference
Although Theorem 1 guarantees efficiency gain under established mild conditions on outcome models, regardless of the choice of variable selection method, the uncertainty introduced by variable selection must be accounted for when conducting inference with the final AIPW estimator. This challenge is commonly referred to as the post-selection inference problem. In this section, we outline the theoretical conditions under which a variable selection method enables valid inference using the variance estimator applied to the AIPW estimator constructed from the selected covariates. Design-based theoretical results related to Lasso-based variable selection, in conjunction with the ANHECOVA estimator, have been studied in Bloniarz et al. (2016).
@@ 3.4.2 Covariate-adaptive randomization
Our second extension centers around the covariate-adaptive randomization (Bugni et al., 2018; Ye et al., 2022; Rafi, 2023; Bannick et al., 2025) including stratified permuted block randomization (Zelen, 1974) and Pocock and Simon’s minimization (Pocock and Simon, 1975; Taves, 1974) as special cases. In general, we let $\mathbf{Z}$ be a discrete baseline covariate vector used in covariate-adaptive randomization, where $\mathbf{Z}$ has $L<\infty$ distinct joint levels $\mathbf{z}_{1},\dots,\mathbf{z}_{L}$. Also, $\mathbf{Z}$ and $\mathbf{X}$ can share some common components. For example, the gender can be a factor for stratified randomization at the design stage, but it can also be incorporated into the covariate adjustment in the analysis stage. Following Bannick et al. (2025), we impose the following mild assumption on treatment assignment in covariate-adaptive RCTs. We use $\pi_{a}(\mathbf{z})=P(A=a\mid\mathbf{Z}=\mathbf{z})$ to denote the treatment assignment probability for arm $a$ within stratum $\mathbf{z}$.
@@
Furthermore, $\mathbf{\Omega}(\mathbf{z})=\text{diag}\left\{\pi_{0}(\mathbf{z}),\pi_{1}(\mathbf{z})\right\}-\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)^{\prime}\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)$ is a $2\times 2$ covariance matrix determined by the covariate-adaptive randomization design. In contrast, $\mathbf{\Omega}_{\text{SR}}=\text{diag}\{\pi_{0},\pi_{1}\}-(\pi_{0}~~\pi_{1})^{\prime}(\pi_{0}~~\pi_{1})$ denotes the corresponding matrix under simple randomization.
@@ 4 Numerical Experiments
First, we generate 20 independent super-population full datasets, each containing $5\times 10^{7}$ independent units, with both potential outcomes $(Y(0),Y(1))$ available for each unit. We then compute the mean of Y(1)Y(0)Y(1)-Y(0) across these 20 large datasets to obtain the true ATE. Given the large sample size and 20 replications, the associated uncertainty is negligible.
@@ 4.3 Performance criteria
  • Power%: The proportion of replications where the p-value is $\leq 0.05$ for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
With $M=500$ simulations, a CP% is considered significantly different from the nominal 95% coverage level if it falls outside the range $95\pm 1.96\times\sqrt{95\times 5/500}=[93,97]$.